Find the thickness of a plate which will produce a change in optical path equal to half the wavelength λ of the light passing through it normally. The refractive index of the plate is μ.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Let, t = thickness of the plate Given, optical path difference = (μ – 1)t = λ/2 ⇒ t = λ/2(μ – 1)